OpenStudy (anonymous):

I'm having a very hard time with partial differentiation find x/y ln(xy+yz+xz)

OpenStudy (anonymous):

please forgive me i dont know how to make the partial derivative sign so thats why it isnt there in the question

OpenStudy (anonymous):

I keep getting -1/2

OpenStudy (anonymous):

its easy ;) you must firstly try refering books then try to solve simple ones and then get to hard ones ;) its an easy topic ;)

OpenStudy (anonymous):

ok, can you help me solve this one. I think I've got it down but this one is giving me trouble

OpenStudy (anonymous):

first of all use chain rule, and if you find two variables at a time ,use product rule note: \[\frac{ d }{ dx }\times y = \frac{ dy }{ dx }\] \[\frac{ d }{ dx } \times y^2 = 2y \times \frac{ dy }{ dx }\] \[\frac{ d }{ dx } xy = x \times \frac{ dy }{ dx } + y \times \frac{ d }{ dx }x\]

OpenStudy (anonymous):

will it be solved by you now?

OpenStudy (dumbcow):

if the function is called say F, then \[\frac{dx}{dy} = \frac{dF}{dy}*\frac{dx}{dF}\] \[\frac{dF}{dy} = \frac{x+z}{xy+xz+yz}\] \[\frac{dF}{dx} = \frac{y+z}{xy+xz+yz}\]

OpenStudy (anonymous):

@dumbcow all your d-s should be written \partial - s

OpenStudy (dumbcow):

*sorry i should have used "partial derivative" notation there

OpenStudy (dumbcow):

the difference is you treat other variables like constants

OpenStudy (turingtest):

@m.auld64 I know that there are many explanations above me, but would you like me to have a crack at it as well?

OpenStudy (anonymous):

sure thing!

OpenStudy (anonymous):

be thorough

OpenStudy (anonymous):

@dumbcow, the notation you have is NOT FOR partial, it can mean different things, it indicates "total derivative" which comprises of partials